# Self-Ensembling Gaussian Splatting for Few-Shot Novel View Synthesis

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS). However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization ability to novel views. In this paper, we alleviate the overfitting problem, presenting a Self-Ensembling Gaussian Splatting (SE-GS) approach. Our method encompasses a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model serves as an ensemble of 3DGS models that generates novel-view images during inference. We achieve the self-ensembling by introducing an uncertainty-aware perturbation strategy at the training state. We complement the $\mathbf{\Sigma}$-model with the $\mathbf{\Delta}$-model, which is dynamically perturbed based on the uncertainties of novel-view renderings across different training steps. The perturbation yields diverse temporal samples in the Gaussian parameter space without additional training costs. The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and these temporal samples. Therefore, our SE-GS conducts an effective and efficient regularization across a large number of 3DGS models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model. Our experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods. The code is released at:project page.
@@ 1 Introduction
Novel view synthesis (NVS) is a critical task [49] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [50], and 3D content generation [16, 42]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [28, 39] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [21, 6, 46] has emerged as a powerful representation, integrating the advantages of both explicit [35] and implicit [28] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints. However, 3DGS tends to overfit the available views when only a limited number of images are provided. To illustrate this issue, in Fig. 2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting problem becomes more noticeable with fewer training views, such as when using only 3 views.
@@
Figure 2: Overfitting in 3D Gaussian Splatting with sparse training views. 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across different training iterations.
To mitigate overfitting, self-ensembling [11, 22, 29] has been highlighted as an effective strategy in tasks such as detection [48] and segmentation [25]. Nevertheless, its use for NVS remains unstudied, and how to exploit it in a 3DGS formalism is an open question. Therefore, in this paper, we bridge this gap, introducing a new 3DGS method that enhances the quality of novel view synthesis with sparse training views via self-ensembling. Our method consists of two models, a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model constitutes an ensemble of 3DGS models, which is used to generate novel views during inference. A straightforward way to achieve ensembling would be to utilize multiple 3DGS models that correspond to different samples in the Gaussian parameter space, thereby increasing the likelihood of reaching a more robust solution. Nevertheless, as we will demonstrate in Sec. 4, the computational cost of training multiple 3DGS models limits the ability to produce sufficiently diverse samples. By contrast, we complement the $\mathbf{\Sigma}$-model with a single $\mathbf{\Delta}$-model and obtain diverse samples by perturbing the $\mathbf{\Delta}$-model. Specifically, instead of applying random perturbation, which often results in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data. We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views at different training iterations in buffers and calculate pixel-level uncertainties across images within each buffer. We then add random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores. Consequently, our perturbation strategy generates uncertainty-aware samples without introducing significant training overhead.
Building upon the uncertainty-aware samples derived from the $\mathbf{\Delta}$-model, we construct an ensemble by leveraging a self-ensembling mechanism. We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbation and regularize the geometry of the $\mathbf{\Sigma}$-model by minimizing its discrepancies with the perturbed $\mathbf{\Delta}$-model. These discrepancies are measured via a photometric loss between images synthesized from the pseudo views. Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving its generalization ability to novel views. Moreover, our self-ensembling technique is independent of any additional ground-truth signals as the regularization is carried out in a self-supervised manner.
@@
We conduct experiments on multiple datasets, including LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45], with limited numbers of training views. Our SE-GS achieves the best performance across all datasets in the sparse-view setting, surpassing the state-of-the-art approaches. Furthermore, we perform a comprehensive analysis of our method to demonstrate its effectiveness and efficiency. To the best of our knowledge, we are the first to explore the potential of the self-ensembling mechanism in 3DGS for few-shot novel view synthesis. The code will be publicly available upon acceptance.
@@ 2 Related Work
Radiance field modeling. Recently, Neural Radiance Field (NeRF) [28, 41, 32, 3] has brought groundbreaking advancements to novel view synthesis. NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9]. However, most NeRF-based methods are inefficient [18] during training and inference because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. In this context, 3D Gaussian Splatting [21] is developed as an efficient alternative for radiance field modeling. 3DGS relies on explicit representations [35] that allow for faster rendering and training [10]. While initially introduced as a scene-specific model, some 3DGS variants [40, 5, 26, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
@@
Ensembling. Ensembling has been evidenced as a powerful technique in machine learning [52, 1, 14] to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [13, 33, 23] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [22] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present the first Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples in the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
@@ 3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into the 3DGS framework. Our self-ensembling builds upon diverse samples in the Gaussian parameter space. These samples are derived from the $\mathbf{\Delta}$-model via an uncertainty-aware perturbation strategy. Specifically, we store images rendered from pseudo views at different training steps in buffers, from which we compute pixel-level uncertainties. We then perturb the Gaussians overlapping the pixels with high uncertainties, as highlighted as red ellipses. Self-ensembling over the resulting samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. During inference, novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
3.1 Preliminaries and Motivations
@@
The parameters of the Gaussians are then learned from the input views. To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, it is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane relies on the projection matrix $\mathbf{W}$ to compute the projected 2D covariance matrix
@@
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion (SFM) techniques [35, 36]. These parameters are optimized via gradient descent with a photometric loss [21] where the posed training images serve as ground truth.
In this paper, we focus on sparse-view scenarios, where only few-shot training images are provided. We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in a suboptimal solution, which ultimately degrades the quality of the synthesized novel views. To handle the overfitting problem, a promising solution is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples. By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space. As we will report in Sec. 4, the performance on testing data improves as $k$ grows. However, training multiple 3DGS models incurs a significant computational cost, which becomes unaffordable with a large $k$. This limitation prevents the method from scaling up for better performance. Consequently, we propose a novel training strategy based on self-ensembling [11, 22], which enhances 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is illustrated in Fig. 3 and the details will be elaborated in the remainder of this section.
@@ 3.2 Uncertainty-Aware Perturbation
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer that stores the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to compute a perturbation that results in a new 3DGS model.
In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training. Note that, in this paper, temporal refers to the dynamic nature of samples that vary across training iterations. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [21]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A temporal 3DGS model is then sampled from the neighboring region of the current $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as
@@
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}$ indicates the perturbation, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model. To achieve the perturbation, a naive approach is to sample random noise from a normal distribution, i.e., $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t}^{2})$. However, such perturbation is independent of the training data, and this method thus tends to yield invalid samples.
@@
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. We then compute a pixel-wise uncertainty map
@@
of the same resolution as the rendered image, i.e., $\mathbf{U}\in\mathbb{R}^{H\times W}$, where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all $M$ sampled pseudo views in parallel. In practice, to enhance robustness, we perform local smoothing over each uncertainty map, which yields
$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(i+m,j+n),$ (9)
where $k=5$ is the kernel size. Subsequently, we apply an uncertainty-aware perturbation to each of the Gaussians in the $\mathbf{\Delta}$-model as
@@
where $\hat{\mathcal{U}}^{t}$ is the set of the uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
@@
with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel sampled from the set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. In short, the reliability of each Gaussian is connected with the uncertainties of the renderings, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed. In practice, we add the perturbation to the 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model. In particular, since 3D rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [51] of the rotation, which yields
@@
where $f(\cdot)$ denotes the mapping from rotation matrix to 6D representation. Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to compute uncertainties directly in the Gaussian parameter space across different training steps. Our approach naturally handles this challenge as the statistics on 2D renderings provide a consistent ground to measure uncertainties, regardless of the varying number of Gaussians.

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness in novel view synthesis (NVS). However, 3DGS tends to overfit when trained with sparse views, limiting its generalization to novel viewpoints. In this paper, we address this overfitting issue by introducing Self-Ensembling Gaussian Splatting (SE-GS). We achieve self-ensembling by incorporating an uncertainty-aware perturbation strategy during training. A $\mathbf{\Delta}$-model and a $\mathbf{\Sigma}$-model are jointly trained on the available images. The $\mathbf{\Delta}$-model is dynamically perturbed based on rendering uncertainty across training steps, generating diverse perturbed models with negligible computational overhead. Discrepancies between the $\mathbf{\Sigma}$-model and these perturbed models are minimized throughout training, forming a robust ensemble of 3DGS models. This ensemble, represented by the $\mathbf{\Sigma}$-model, is then used to generate novel-view images during inference. Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets demonstrate that our approach enhances NVS quality under few-shot training conditions, outperforming existing state-of-the-art methods. The code is released at: project page.
@@ 1 Introduction
Novel view synthesis (NVS) is a critical task [50] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [51], and 3D content generation [16, 43]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [29, 40] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [21, 6, 47] has emerged as a powerful representation, integrating the advantages of both explicit [36] and implicit [29] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints. However, 3DGS tends to overfit the available views when only a limited number of images are provided. To illustrate this issue, in Fig. 2, we evaluate 3DGS trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting problem becomes more noticeable with fewer training views, such as when using only 3 views.
@@
Figure 2: Overfitting in 3D Gaussian Splatting with sparse training views. (a) and (b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across training iterations.
To mitigate overfitting, ensembling [23, 22, 10] has been highlighted as an effective strategy in detection [49] and segmentation [26]. Nevertheless, its use for NVS remains unstudied, and how to exploit it in a 3DGS formalism is an open question. Therefore, in this paper, we bridge this gap, introducing a new 3DGS method that enhances the quality of novel view synthesis with sparse training views via self-ensembling. A straightforward way to achieve ensembling would be to train multiple 3DGS models and combine the corresponding predictions as the final result during testing. However, as we will demonstrate in Sec. 4, this method is computationally expensive, and the resulting models lack sufficient diversity to ensure effective ensembling.
In contrast, we present an effective and efficient self-ensembling mechanism. In ensemble learning [11, 30], self-ensembling is typically achieved by introducing perturbations during training through augmentation or dropout. Inspired by its success, we introduce an uncertainty-aware perturbation strategy for 3DGS. Specifically, we train a $\mathbf{\Delta}$-model on available RGB images. During training, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the renderings. To compute these uncertainties, we store images rendered from the $\mathbf{\Delta}$-model at pseudo views across different training iterations in buffers and calculate pixel-level uncertainties within each buffer. A perturbed model is then obtained by adding random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores. Since all perturbed models are derived from the $\mathbf{\Delta}$-model rather than being trained from scratch, our perturbation strategy generates diverse models without incurring significant computational overhead. In addition to the $\mathbf{\Delta}$-model, we train a $\mathbf{\Sigma}$-model on the training views without the perturbation. The self-ensembling is achieved by minimizing the discrepancies between the $\mathbf{\Sigma}$-model and perturbed models. These discrepancies are measured via a photometric loss between images synthesized at the pseudo views. This self-ensembling is performed based on diverse models, thereby improving the robustness and generalization of the resulting ensemble, the $\mathbf{\Sigma}$-model. Moreover, our self-ensembling is independent of any additional ground-truth signals as it is carried out in a self-supervised manner. During testing, the $\mathbf{\Sigma}$-model is employed for novel view synthesis.
@@
We conduct experiments on multiple datasets, including LLFF [28], DTU [20], Mip-NeRF360 [3], and MVImgNet [46], with sparse training views. Our SE-GS achieves the best performance across all datasets in the sparse-view setting, surpassing the state-of-the-art approaches. Furthermore, we perform a comprehensive analysis of our method to demonstrate its effectiveness and efficiency. To the best of our knowledge, we are the first to explore the potential of the self-ensembling mechanism in 3DGS for few-shot novel view synthesis. The code will be publicly available upon acceptance.
@@ 2 Related Work
Radiance field modeling. Recently, Neural Radiance Field (NeRF) [29, 42, 33, 3] has brought groundbreaking advancements to novel view synthesis. However, most NeRF-based methods are inefficient [18] because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. In this context, 3D Gaussian Splatting [21] is developed as an efficient alternative for radiance field modeling. 3DGS relies on explicit representations [36] that allow for faster rendering and training [9]. While initially introduced as a scene-specific model, some 3DGS variants [41, 5, 27, 7] explore the generalization of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
@@
Ensemble learning. Ensembling has been evidenced as a powerful technique in machine learning [53, 1, 14] to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [13, 34, 24] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [22] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present the first Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples in the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
@@ 3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into 3DGS. We jointly train a $\mathbf{\Delta}$-model and a $\mathbf{\Sigma}$-model. During training, we store pseudo-view renderings of the $\mathbf{\Delta}$-model in buffers, from which we compute pixel-level uncertainties. The Gaussians of the $\mathbf{\Delta}$-model overlapping the pixels with high uncertainties are perturbed, as highlighted as red ellipses, which leads to a perturbed model. We then achieve self-ensembling by penalizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed models. During inference, the resulting ensemble, the $\mathbf{\Sigma}$-model, is used for novel view synthesis.
3.1 Preliminaries
@@
To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, it is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane relies on the projection matrix $\mathbf{W}$ to compute the projected 2D covariance matrix
@@
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion techniques [36, 37]. These parameters are optimized with a photometric loss [21] where the posed training images serve as ground truth.
3.2 Motivations
In this paper, we focus on sparse-view scenarios, where few-shot images are provided. We denote the 3DGS model trained on these images as $\mathcal{G}$. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in a suboptimal solution, which ultimately degrades the quality of the synthesized novel views. A promising approach to mitigate overfitting is ensemble learning [13], which has been utilized to enhance robustness and generalization in the literature. Rather than depending on a single model, ensemble learning aggregates predictions from multiple models, thereby stabilizing the final output. To achieve this, a straightforward method is to jointly train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$. As we will report in Sec. 4, by aggregating their predictions, the final results are more robust, and the NVS performance improves as $k$ grows. However, training multiple 3DGS models incurs significant computational costs, making it impractical for large $k$. Moreover, our experimental findings indicate that the trained 3DGS models are not diverse enough to support effective ensembling. Consequently, we propose a novel self-ensembling paradigm based on an uncertainty-aware perturbation strategy, enhancing 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is shown in Fig. 3 and the details will be elaborated in this section.
@@ 3.3 Uncertainty-Aware Perturbation
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer storing the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to determine perturbation that results in a new 3DGS model.
In contrast to training separate 3DGS models from scratch, we dynamically generate diverse models from a single $\mathbf{\Delta}$-model during training. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [21]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A new 3DGS model is then created by perturbing the $\mathbf{\Delta}$-model as
@@
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}\in\mathcal{N}(\mathbf{\mu_{t}},\mathbf{\sigma}^{2}_{t})$ indicates the noise, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the perturbed model. This naive approach adds random noise to all Gaussians in the $\mathbf{\Delta}$-model. However, as we will demonstrate in Sec. 4, this method shifts perturbed models too far from the $\mathbf{\Delta}$-model, leading to instability and unreliable supervision.
@@
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a randomly sampled scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. We then compute a pixel-wise uncertainty map
@@
where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all $M$ sampled pseudo views in parallel. To enhance robustness, we perform local smoothing over each uncertainty map, yielding
$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(m,n),$ (9)
where $k=5$ is the size of a kernel centered at $(i,j)$ and $(m,n)$ indicate the coordinates of a pixel within the kernel. Subsequently, we apply uncertainty-aware perturbation to each Gaussian in the $\mathbf{\Delta}$-model as
@@
where $\hat{\mathcal{U}}^{t}=\{\hat{\mathbf{U}}^{t}_{1},\hat{\mathbf{U}}^{t}_{2},\cdots,\hat{\mathbf{U}}^{t}_{M}\}$ is a set of uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
@@
with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel in a set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. In short, the reliability of each Gaussian is connected with the rendering uncertainty, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed. In practice, we add random noise to 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model. Particularly, since rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [52], which is denoted as
@@
where $f(\cdot)$ indicates the mapping from rotation matrix to 6D representation. Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse perturbed models from the $\mathbf{\Delta}$-model.
It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to assess uncertainties directly in the Gaussian parameter space for newly generated Gaussians. Our approach naturally handles this challenge as the statistics on 2D renderings provide a consistent ground to measure uncertainties, regardless of the varying number of Gaussians.